# The Transfer Performance of Economic Models

### v4

@@ 1. Introduction
Economic models estimated on data from one context are often used to guide predictions and policy in a range of other contexts. For example, a model of information diffusion estimated on data of microfinance takeup in one Indian village may be used to guide policy decisions about the seeding of microfinance in another village, and a model of risk preferences estimated on willingness-to-pay data for certain lotteries may be used to predict willingness-to-pay for new lotteries. How can the generalizability of a model to new settings be assessed and predicted?
The generalizability of models is a classic concern in economics (Haavelmo, 1944; Chassang and Kapon, 2022; DellaVigna and Pope, 2019), but has a new importance due to the increased popularity of “black-box” machine learning prediction methods. Machine learning methods have been shown to out-predict economic models (Hartford et al., 2016; Plonsky et al., 2019; Hofman et al., 2021) and identify new interpretable regularities that existing models do not capture (Fudenberg and Liang, 2019; Peterson et al., 2021; Ludwig and Mullainathan, 2023). At the same time, many believe that structured economic models capture regularities that generalize well across domains, and are more reliable for making predictions in new contexts (Coveney et al., 2016; Athey, 2017; Beery et al., 2018; Manski, 2021). Whether economic models in fact generalize better is an empirical question.
Our paper’s contribution is twofold. First, we provide a tractable approach for evaluating cross-domain transfer performance based on techniques that generalize conformal inference (e.g. Vovk et al., 2005; Barber et al., 2021; Angelopoulos et al., 2022). In our statistical model, behavior in different economic contexts is governed by different distributions, which are themselves drawn i.i.d. from a fixed but unknown meta distribution. [^1] Section 3.2 and Appendix R.1 relax the i.i.d. assumption. Under this assumption, we derive finite-sample forecast intervals for a large class of measures of transfer performance; these intervals can be used to evaluate economic models, regression models, and black box algorithms alike.222 We use the term “forecast interval,” rather than “confidence interval,” to reflect the random nature of the target, namely the realized (rather than expected, median, etc.) transfer error, but they can also be viewed as confidence intervals for these random targets. Second, we use these forecast intervals to compare the generalizability of economic models and black box machine learning methods in a specific economic application (predicting certainty equivalents for lotteries) and find that economic models generalize better.
@@
Our conceptual framework, described in Section 2, is an extension of the familiar notion of “out-of-sample” evaluation to “out-of-domain” evaluation. In the standard out-of-sample test, a model’s free parameters are estimated on a training sample, and the predictions of the estimated model are evaluated on a test sample, where the observations in the training and test samples are disjoint but drawn from the same distribution. We depart from this framework by supposing that the distribution of the data varies across a set of “domains,” where these domain-specific distributions are drawn i.i.d. from a meta-distribution. As Section 2 explains, our results apply to a large class of measures for the transferability for a model, which we call transfer errors. Transfer errors can be used to evaluate the performance of many common empirical techniques, including using a model that is trained on a sample from one domain to predict in a sample from an as-yet unobserved domain, and asking whether a qualitative prediction based on estimated parameters from one sample will generalize to another.
Section 3 shows how to construct forecast intervals with guaranteed coverage probability for any transfer error, using a meta-data set of samples from already observed domains. Our approach is to split the observed domains in the meta-data set into training and test domains, estimate the parameters of the model on the samples from the training domains, and evaluate its transfer error on each of the test domains. Pooling these transfer errors across different choices of training and test domains yields an empirical distribution of transfer errors. We show that for every quantile $\tau$, the interval bounded by the $\tau$-th and $(1-\tau)$-th quantiles of the pooled transfer error is a valid forecast interval for the transfer error on a new, unseen domain (up to a finite-sample correction). We also relax the i.i.d. sampling assumption, deriving a modified procedure for cases where the distributions in training domains are drawn i.i.d. from one distribution, while the distribution in the target domain is drawn from another. Both procedures (and all other methods described in this paper) are implemented in an R package (transferUQ), available on Github. [^3] https://github.com/lihualei71/transferUQ
@@
Section 4 applies these procedures to compare the transferability of economic models and black box algorithms in a classic economic problem: predicting certainty equivalents for binary lotteries. The samples correspond to observations from different subject pools, so a model’s transfer error describes how well it predicts outcomes in one subject pool when estimated on data from another. We evaluate two models of risk preferences, expected utility and cumulative prospect theory, and two popular black box machine learning algorithms, random forest and kernel regression. We find that although the black box algorithms outperform the economic models out-of-sample when trained and tested on data from the same domain, the economic models generalize more reliably across domains. Specifically, while the forecast intervals for the black box algorithms and economic models overlap, the forecast intervals for the black box methods are wider, and their upper bounds are substantially higher.
@@
Why do the black boxes perform worse at transfer prediction in this setting? A natural explanation, based on intuition from conventional out-of-sample testing, is that black boxes are very flexible and hence learn idiosyncratic details that do not generalize across subject pools. But when we restrict the analysis to a subset of samples involving the same set of lotteries, the resulting forecast intervals are nearly identical across all of the prediction methods. This tells us that black box methods do not always transfer worse. Instead, black boxes seem to transfer worse when the primary source of variation across samples is a shift in the marginal distribution over features (i.e., which lotteries appear in the sample), rather than a shift in the distribution of outcomes conditional on features (i.e., the distribution of certainty equivalents given fixed lotteries). We find further evidence for this when we consider an alternative definition of the domains, where the transfer task corresponds to training on some lotteries and predicting for others. Here the improvement of the transfer performance of economic models over black box algorithms is even larger than in our main analysis. Taken together, these results suggest that the crucial difference between economic models of risk preferences and black box algorithms is not that one is more flexible than the other, but rather that economic models do a better job of relating behavior across lotteries.
@@
Although our results apply for a broad class of definitions of transfer error, our primary motivation is evaluating how well an economic model estimated on data from one domain predicts in another. Hofman et al. (2021) gives an in-depth argument for why this important, calling for more work on the question “how well does a predictive model fit to one data distribution generalize to another?” for social science models. This is exactly what we consider.
@@
Predictive accuracy is not the only criterion that matters for evaluating models, but it has a central role in experimental economics. As discussed by Harless and Camerer (1994), the poor predictive performance of expected utility theory was a primary motivation for the development of alternative models in behavioral economics, and both Harless and Camerer (1994) and Hey and Orme (1994) provide early assessments of alternative theories on the basis of predictive performance. More recently, several papers have evaluated how well the predictions of economic models transfer across domains. For example, Külpmann and Kuzmics (2022) estimates various game-theoretic models on $2\times 2$ normal-form games and evaluates their predictive performance on $3\times 3$ normal-form games; Natenzon (2019) estimates discrete choice models on data for four choice menus and evaluates their predictive performance on a fifth menu; and Fudenberg and Karreskog Rehbinder (2024) evaluates the cross-domain predictive accuracy of models of how players learn in the infinitely-repeated prisoner’s dilemma, where the domains are various sets of payoff matrices for the stage game. This paper provides a general framework that nests these transfer problems, and formal statistical results for assessing transfer performance.
@@
Our comparison of economic models and black box algorithms joins a small but growing body of work comparing the value of these methods (Athey and Imbens, 2016; Fudenberg and Liang, 2019; Agrawal et al., 2020). Several recently published papers compare the predictive performance of black box algorithms with that of more structured economic models in out-of-sample tests, where the training and testing data are drawn from the same domain (Peysakhovich and Naecker, 2017; Noti et al., 2016; Plonsky et al., 2019; Camerer et al., 2019; Fudenberg and Liang, 2019; Hirasawa et al., 2022; Hsieh et al., 2023). In contrast, our paper compares economic models and black box algorithms from the perspective of transfer prediction. We find that although black box methods are often very effective given a large quantity of data from the domain of interest, they may be less effective at transferring predictions across domains. [^4] At a high level, this is similar to Gechter et al. (2019)’s finding that structural models deliver better policy recommendations for new contexts than black box methods do.
@@
Finally, our theoretical framework and results lie at the intersection of several literatures in economics, computer science, and statistics. These literatures consider several related but distinct tasks: synthesizing evidence across different domains, improving the quality of extrapolation from one domain to another, and quantifying the extent to which insights from one domain generalize to another. Our results are most closely related to this third strand.
@@
The first objective, synthesizing results across different domains, is a particular focus of the literature on meta-analysis (Card and Krueger, 1995; DellaVigna and Pope, 2019; Bandiera et al., 2021; Imai et al., 2020; Vivalt, 2020). Our goal is instead to assess the cross-domain forecast accuracy of a model. These problems are related, and Meager (2019) and Meager (2022) in particular provide posterior predictive intervals for new domains in the context they consider. Unlike our approach, the predictive intervals reported in those papers are valid only under a parametric model for the distribution of effects across domains.
@@
There is also a large literature that aims to extrapolate results from one domain to another. Within computer science, the literature on domain generalization (Blanchard et al. 2011 and Muandet et al. 2013) develops models that generalize well to new unseen domains (Zhou et al., 2021). Similarly, several papers within economics (e.g., Hotz et al. 2005 and Dehejia et al. 2021) use knowledge about the distribution of covariates to extrapolate out-of-domain. Our focus is not on developing new models or algorithms with good out-of-domain guarantees, but rather on developing forecast intervals for the out-of-domain performance of models and algorithms that are used in practice.
@@
Finally, the literature on external validity studies the extent to which results obtained in one domain hold more generally. This paper does not focus on the generalizability of insights from randomized control trials (e.g. Deaton, 2010; Imbens, 2010; DellaVigna and Linos, 2020) or laboratory experiments (e.g. Levitt and List, 2007; Al-Ubaydli and List, 2015), but instead on a model’s generalizability across exchangeable domains. [^5] Another set of papers study the generalizability of instrumental variables estimates (e.g. Angrist and Fernández-Val, 2013; Bertanha and Imbens, 2020) and causal effects (e.g. Pearl and Bareinboim, 2014; Park et al., 2023). Our use of exchangeability to construct bounds extends work on conformal inference (e.g. Vovk et al., 2005; Barber et al., 2021; Angelopoulos et al., 2022) by replacing the assumption of exchangeable observations with that of exchangeable domains. [^6] This also differentiates our work from the out-of-distribution prediction literature in computer science (Shen et al., 2021), which bounds expected transfer error when the test and training distributions are close. Section 3.2 relaxes this assumption; our results there connect to the literature on sensitivity analysis (e.g. Rosenbaum, 2005; Aronow and Lee, 2013; Andrews and Oster, 2019).
@@ 2. Framework
Consider a fixed procedure for extrapolating predictions across domains, e.g., estimating a structural economic model on data from one domain and using the estimated model to make predictions in another. We adopt the perspective of an external analyst who wants to evaluate the effectiveness of this procedure, using data from multiple domains. The analyst is not focused on extrapolation from one specific domain to another (e.g., from an American dataset to a German dataset), but would rather like to understand whether the procedure generally performs well across a class of transfer tasks (e.g., extrapolating across countries). To this end, the analyst evaluates transfer error from an ex-ante perspective without knowing which domains are used to estimate and evaluate the model. The analyst seeks to construct forecast intervals for the procedure’s error when transferring from a (random) set of training domains to a new (random) target domain. This speaks to the question of whether one procedure for extrapolation (such as transferring an estimated economic model) generally performs better than another (such as transferring a trained black box algorithm).
This section proceeds as follows: Section 2.1 formalizes the extrapolation procedure that the analyst would like to evaluate, and a large class of measures for the procedure’s transfer error. Section 2.2 describes the analyst’s problem, and in particular the statistical assumption that permits the analyst to construct forecast intervals for these transfer errors. Section 2.3 defines these intervals, and Section 3 proves that these are in fact valid forecast intervals.

### v5

@@ 1. Introduction
Economists routinely make predictions in environments where data is unavailable, relying on evidence from related but distinct contexts. For example, an economist at a development agency may need to predict diffusion of microfinance takeup in one Indian village given data on diffusion in others. Or an economist at an insurance company may need to predict willingness-to-pay for certain insurance plans given data on willingness-to-pay for others.
Traditionally, economists address this challenge by estimating economic models on existing data and using the estimated parameters to make predictions in new domains. But the predictive success of machine learning algorithms in several economics applications (e.g. Hartford et al., 2016; Hofman et al., 2021; Banerjee et al., 2024) raises the question of whether the economic model is essential in this process, or whether predictions would be improved by training and porting a flexible machine learning algorithm instead. This is an empirical question: On the one hand, machine learning methods are capable of uncovering novel patterns that existing models miss (Fudenberg and Liang, 2019; Peterson et al., 2021; Ludwig and Mullainathan, 2023); on the other, many researchers believe that structured economic models capture fundamental regularities that generalize more reliably across domains (Coveney et al., 2016; Athey, 2017; Beery et al., 2018; Manski, 2021). Understanding whether economic models do in fact transfer better across domains is important to understanding their future role within economic analysis and policymaking.
Our paper provides a conceptual framework that formalizes the “out-of-domain” transfer problem, and allows systematic comparison between economic modeling and machine learning algorithms. Relative to the large body of work on external validity described in Section 1.1, our out-of-domain test considers performance within a class of prediction problems rather than conditioning on a particular realization of the data available to the economist. For example, instead of assessing whether a structural model of information diffusion will transfer well from a specific Indian village to another, we ask whether the structural model transfers well “in general” across Indian villages. Formally, we consider an ex-ante perspective in which the economist is aware of a set of domains that are relevant (such as Indian villages) but does not know which specific transfer problem will realize from this class of possible transfer problems.
Our main theoretical contribution is the construction of finite-sample forecast intervals that characterize how well a model or algorithm performs in new domains.111 We use the term “forecast interval,” rather than “confidence interval,” to reflect the random nature of the target, namely the realized (rather than expected, median, etc.) transfer error, but they can also be viewed as confidence intervals for these random targets. In the main text, we provide intervals that are valid for a benchmark setting where all domains are equally likely to realize for training and testing (corresponding to an assumption that domains are exchangeable) as well as for a setting in which the testing domain is qualitatively different from the training domains. (Appendix R.1 generalizes our approach and results even further.) The definition of “performance” is left to the user, and our results apply for a large class of measures that includes our motivating example of predictive accuracy, as well as others, such as how well a qualitative finding based on parameter estimates generalizes across domains.
Finally, we use our framework to compare the generalizability of economic models and black box machine learning methods when predicting certainty equivalents for lotteries. While economic models of risk preference have been studied extensively from the perspectives of how well they fit economic data (Harless and Camerer, 1994; Hey and Orme, 1994; Bruhin et al., 2010; Bernheim and Sprenger, 2020) and how their predictive performance compares to that of machine learning algorithms (Peysakhovich and Naecker, 2017; Plonsky et al., 2019; Fudenberg et al., 2022), the existing tests have primarily been within the context of a single domain. [^2] An important exception is Einav et al. (2012), which examines how general risk preferences are across different domains of choice under uncertainty. We offer a new perspective on how well these models transfer across domains, finding that although machine learning algorithms outperform the economic models when trained and tested on (disjoint) data generated under the same conditions, the economic models generalize better. Our analysis suggests the primary reason for this is not that the machine learning algorithms overfit, but rather that economic models of risk preference are more effective at extrapolating from observed cases to new ones.
@@
We now describe the main parts of the paper in more detail. Section 2 describes our conceptual framework, which extends the familiar notion of out-of-sample evaluation to out-of-domain evaluation. In the standard out-of-sample test, a model’s free parameters are estimated on a training sample, and the predictions of the estimated model are evaluated on a test sample, where the observations in the training and test samples are disjoint but drawn from the same distribution. We depart from this framework by supposing that the distribution of the data varies across a set of “domains,” e.g. different subject pools or choice frames.
We adopt the perspective of an external analyst who wants to assess the efficacy of a procedure that a researcher uses to make predictions in a new domain. For example, the researcher’s procedure might be to estimate an economic model on data from one domain and use the estimated model to make predictions in another. Alternatively, the researcher’s procedure might involve pooling data across domains for estimation, or instead training and porting a black box algorithm. The procedure’s exact performance depends on the domains that are used for estimation—the training domains—and the domain that is realized for prediction—the target domain. The analyst’s goal is to develop a forecast interval for the random performance of this procedure across the many possible realizations of those domains. This forecast interval can then be used to compare procedures for generalization.
Section 3 constructs forecast intervals for the baseline setting where the distributions governing different domains are exchangeable. This means that while there may be ex-post differences between the domains on which the model is estimated and on which it is applied, these differences are not ex-ante known to the analyst. We propose the following protocol: The analyst first collects data across as many domains as possible, henceforth the analyst’s meta-data set. The analyst then repeatedly splits these into training and test domains, and runs the researcher’s procedure for each of these domain realizations. For example, if the researcher’s procedure involves transferring an economic model, then the analyst would estimate the researcher’s model on the training domains and use it to make predictions in the test domain. Intuitively, the transfer performance between domains in the analyst’s meta-data set can serve as a proxy for the transfer performance to the new (unobserved) target domain. Pooling these transfer performances across different choices of training and test domains yields an empirical distribution of transfer errors. We show that the quantiles of this distribution can be used to compute valid finite-sample forecast intervals for the transfer error on the target domain. (Because in most applications, models are estimated on data from a small number of domains, we exclusively consider finite-sample results in this paper.)
Section 4 extends our methods for the scenario where the training domains are known to be systematically different from the target domain. Specifically, we suppose that the training domains are exchangeable, but allow the target domain to be governed by a qualitatively different distribution. We fully generalize our procedure and results when the analyst either knows or can bound a likelihood ratio relating the target and training distributions. We further develop two partial orders for comparing how well models generalize in environments where the analyst lacks any knowledge at all about the likelihood ratio. Though this ordering is inherently incomplete, we demonstrate that it can be used to compare economic models and black box algorithms in our subsequent application. Each of these procedures (and all other methods described in this paper) are implemented in an R package (transferUQ), available on Github. [^3] https://github.com/lihualei71/transferUQ
Our statistical framework and results extend the recent literature on conformal inference (e.g. Vovk et al., 2005; Tibshirani et al., 2019; Lei and Candès, 2021) and randomization inference (e.g. Ritzwoller et al., 2024) by replacing the assumption of (weighted) exchangeable observations with that of (weighted) exchangeable domains. Importantly, while conformal inference aims to predict a single data point from other data points, our goal is to generate forecast intervals on a function of multiple data points, involving both observed and unobserved data. Our problem thus involves a multi-dimensional structure that falls outside of the standard framework and requires substantively new arguments.
@@
Section 5 applies our procedures to compare the transferability of economic models and black box algorithms for predicting certainty equivalents for binary lotteries. In this application, we evaluate how well a model (or algorithm) predicts certainty equivalents reported by one subject pool when estimated on data from another. We evaluate two models of risk preferences (expected utility and cumulative prospect theory) and two popular black box machine learning algorithms (random forest and kernel regression). We find that although the black box algorithms outperform the economic models out-of-sample when trained and tested on data from the same domain, the economic models generalize more reliably across domains. Specifically, while the forecast intervals for the black box algorithms and economic models overlap, the forecast intervals for the black box methods are wider, and their upper bounds are substantially higher.
@@
Why do the black boxes perform worse at transfer prediction? A natural explanation, based on intuition from conventional out-of-sample testing, is that black boxes are very flexible and hence learn idiosyncratic details that do not generalize across subject pools. But when we restrict the analysis to a subset of samples involving the same set of lotteries, the economic models and black box algorithms have nearly indistinguishable forecast intervals. Alternatively, when we define the domains so that they involve disjoint lotteries, the improvement of the transfer performance of economic models over black box algorithms is even larger than in our main analysis. This tells us that black box algorithms are not universally worse at transfer, but rather their relative performance depends on specific characteristics of the transfer problem. In particular, black boxes seem to transfer worse when the primary source of variation across samples is a shift in the marginal distribution over features (i.e., which lotteries appear in the sample), rather than a shift in the distribution of outcomes conditional on features (i.e., the distribution of certainty equivalents given fixed lotteries).
@@
Our results apply for a broad class of definitions of transfer error, but our primary motivation is evaluating how well an economic model estimated on data from one domain predicts in another. Hofman et al. (2021) gives an in-depth argument for why this important, calling for more work on the question “how well does a predictive model fit to one data distribution generalize to another?” for social science models. This is exactly what we consider.
@@
Predictive accuracy has long had a central role in experimental economics. [^4] As discussed by Harless and Camerer (1994), the poor predictive performance of expected utility theory was a primary motivation for the development of alternative models in behavioral economics. Both Harless and Camerer (1994) and Hey and Orme (1994) provide early assessments of alternative theories on the basis of predictive performance. While models have often been assessed based on how well they fit data from a given domain, recent papers also examine how well predictions transfer across domains. For example, Külpmann and Kuzmics (2022) estimates various game-theoretic models on $2\times 2$ normal-form games and evaluates their predictive performance on $3\times 3$ normal-form games; Natenzon (2019) estimates discrete choice models on data for four choice menus and evaluates their predictive performance on a fifth menu; and Fudenberg and Karreskog Rehbinder (2024) evaluates the cross-domain predictive accuracy of models of how players learn in the infinitely-repeated prisoner’s dilemma, where the domains are various sets of payoff matrices for the stage game. This paper provides a general framework that nests these transfer problems, and develops formal statistical results for assessing transfer performance.
@@
Transfer prediction is also an important consideration when comparing economic models with black box algorithms. Several papers have compared the predictive performance of machine learning algorithms with that of more structured economic models in out-of-sample tests (Peysakhovich and Naecker, 2017; Noti et al., 2016; Plonsky et al., 2019; Camerer et al., 2019; Fudenberg and Liang, 2019; Hirasawa et al., 2022; Hsieh et al., 2023), concluding that machine learning algorithms are more predictive. In the application we consider, black box methods turn out to be less effective at transferring predictions across domains. This result joins Gechter et al. (2019)’s finding that structural models deliver better policy recommendations for conditional cash transfer policies in new contexts than black box methods do. We hope that our methods will be used to assess the transferability of economic models in other problems as well, contributing to a more comprehensive picture of how well economic models generalize.
@@
Finally, our theoretical framework and results lie at the intersection of several literatures in economics, computer science, and statistics. These literatures consider several related but distinct objectives: synthesizing evidence across different domains, improving the quality of extrapolation from one domain to another, and quantifying the extent to which insights from one domain generalize to another.
@@
The first objective, synthesizing results across different domains, is a particular focus of the literature on meta-analysis (Card and Krueger, 1995; DellaVigna and Pope, 2019; Bandiera et al., 2021; Imai et al., 2020; Vivalt, 2020). Our goal is instead to assess the cross-domain forecast accuracy of a model. These problems are related, and Meager (2019) and Meager (2022) in particular provide posterior predictive intervals for new domains in the context they consider. Unlike our approach, the predictive intervals reported in those papers are valid only under a parametric model for the distribution of effects across domains and a distributional assumption on the domain-specific effect estimates.
@@
The second objective is to extrapolate results from one domain to another. Within computer science, the literature on domain generalization (Blanchard et al. 2011 and Muandet et al. 2013) develops models that generalize well to new unseen domains (Zhou et al., 2021). Similarly, several papers within economics (e.g., Hotz et al. 2005 and Dehejia et al. 2021) use knowledge about the distribution of covariates to extrapolate out-of-domain. Our focus is not on developing new models or algorithms with good out-of-domain guarantees, but rather on developing forecast intervals for the out-of-domain performance of models and algorithms that are used in practice.
@@
The third related objective in the literature is to understand the extent to which results obtained in one domain hold more generally, i.e. their external validity. Unlike papers that assess the generalizability of insights from randomized control trials (e.g. Deaton, 2010; Imbens, 2010; DellaVigna and Linos, 2020) or laboratory experiments (e.g. Levitt and List, 2007; Al-Ubaydli and List, 2015), here we consider generalizability across random domains. [^5] Another set of papers study the generalizability of instrumental variables estimates (e.g. Angrist and Fernández-Val, 2013; Bertanha and Imbens, 2020) and causal effects (e.g. Pearl and Bareinboim, 2014; Park et al., 2023). Section 4 and Appendix R.1 relaxes the exchangeability assumption to allow bounded deviations from exchangeability; our results there connect to the literature on sensitivity analysis (e.g. Rosenbaum, 2005; Aronow and Lee, 2013; Andrews and Oster, 2019; Sahoo et al., 2022).
@@ 2. Framework
Consider a fixed procedure for extrapolating predictions across domains, e.g., estimating a structural economic model on data from one domain and using the estimated model to make predictions in another. We adopt the perspective of an external analyst who wants to evaluate the effectiveness of this procedure. The analyst is not focused on extrapolation from one specific domain to another (e.g., from an American dataset to a German dataset), but would rather like to understand whether the procedure generally performs well across a class of transfer tasks (e.g., extrapolating across countries). To this end, the analyst evaluates transfer error from an ex-ante perspective without knowing which domains are used to estimate and evaluate the model. The analyst seeks to construct forecast intervals for the procedure’s error when transferring from a (random) set of training domains to a new (random) target domain. This speaks to the question of whether one procedure for extrapolation (such as transferring an estimated economic model) generally performs better than another (such as transferring a trained black box algorithm).
This section proceeds as follows: Section 2.1 describes the extrapolation procedure that the analyst would like to evaluate, and a large class of measures for the procedure’s transfer error. Section 2.2 formalizes the analyst’s problem.